Consider the following reaction: Cr2O72– + 6Br– + 14 H+ à 2 Cr3+ + 3Br2 + 7H2O
In a redox titration, 14.62 mL of 0.175 M Cr2O72– was needed to completely react with a 25.00 mL sample of Br–. Calculate the [Br–] in the original solution.
1
Expert's answer
2020-05-24T03:10:34-0400

In the reaction

Cr2O72– + 6Br + 14 H+ "\rightarrow" 2 Cr3+ + 3Br2 + 7H2O,

1 ion of Cr2O72– reacts with 6 ions of Br-, therefore their number of the moles relate as:

"n(Cr_2O_7^{2–}) = \frac{n(Br^-)}{6}" .

The number of the moles of the bichromate anion is:

"n(Cr_2O_7^{2\u2013}) = cV = 14.62\u00b710^{-3}\u00b70.175 = 2.559\u00b710^{-3}" mol.

The number of the moles in 25.000 mL sample of Br- is:

"n(Br^-) = 6\u00b72.559\u00b710^{-3} = 0.0154" mol.

Finally, the concentration of Br- in the original solution is the ratio of the number of the moles of Br- that react in titration to the volume of the sample:

"c(Br^-) = \frac{n}{V} = \frac{0.0154}{25·10^{-3}} = 0.614" M.

Answer: the Br- concentration in the original solution is 0.614 M.